# Fast Policy Learning for Linear Quadratic Regulator with Entropy Regularization

### v1

@@ Appendix B Proofs in Section 3
$=\operatorname{Tr}\left(-\nabla_{\Sigma}C(K,\Sigma)\cdot\left(\frac{\tau}{2}(R+\gamma B^{\top}P_{K}B)^{-1}-\Sigma\right)\right)$
@@
$\overset{(b)}{\leq}(1-\gamma)\|\Sigma\|\cdot\|(R+\gamma B^{\top}P_{K}B)^{-1}\|\cdot\|\nabla_{\Sigma}C(K,\Sigma)\|_{F}^{2}\leq\frac{(1-\gamma)\|\nabla_{\Sigma}C(K,\Sigma)\|_{F}^{2}}{\sigma_{\min}(R)},$
where $(a)$ follows from the first order concavity condition for $f_{K}$ and $(b)$ follows from $\Sigma\preceq I.$
@@
B.3 Proof of Lemma 3.3
Lemma B.2 shows that the cost objective is smooth in $\Sigma$ when utilizing entropy regularization, given that $\Sigma$ is bounded.
Lemma B.2 (Smoothness of $f_{K}$ (16)).
@@
(of Lemma B.2). Fix symmetric positive definite matrices $X$ and $Y$ satisfying $aI\preceq X\preceq I$ and $aI\preceq Y\preceq I$. Then $f_{K}$ being concave implies $f_{K}(X)-f_{K}(Y)+\operatorname{Tr}\left(\nabla f_{K}(X)^{\top}(Y-X)\right)\geq 0.$ To find an upper bound, observe that
@@
(of Lemma 3.3). The first equality immediately results from Lemma B.1, by taking an expectation with respect to $x$ in (46). The last inequality follows directly from Lemma B.2. ∎
Appendix C Proofs in Section 4
C.1 Proofs of Lemma 4.1
@@ Appendix C Proofs in Section 4
C.2 Proof of Lemma 4.2
For ease of exposition, write $S=S_{K,\Sigma},$ and $S^{\prime}=S_{K^{\prime},\Sigma^{\prime}}$. Let $M_{a}$ be defined in the same way as in Lemma 3.3. Let $f_{K}$ be defined as (16). Then Lemma 3.3 implies
$C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)$ $=\operatorname{Tr}\left(S^{\prime}(K^{\prime}-K)^{\top}(R+\gamma B^{\top}P_{K}B)(K^{\prime}-K)\right)$
@@
$\leq$ $4\eta_{1}^{2}\|R+\gamma B^{\top}P_{K}B\|\operatorname{Tr}\left(S^{\prime}E_{K}^{\top}E_{K}\right)-4\eta_{1}\operatorname{Tr}\left(S^{\prime}E_{K}^{\top}E_{K}\right)$
$\overset{(a)}{\leq}$ $-2\eta_{1}\operatorname{Tr}\left(S^{\prime}E_{K}^{\top}E_{K}\right)\leq-2\eta_{1}\mu\operatorname{Tr}\left(E_{K}^{\top}E_{K}\right)\leq-2\eta_{1}\mu\frac{\sigma_{\min}(R)}{\|S^{*}\|}f_{1}(K,K^{*},\Sigma,\Sigma^{*}),$
where the inequality (a) in (54) follows from $\eta_{1}\leq\frac{1}{2\|R+\gamma B^{\top}P_{K}B\|}$.
By Lemma 4.1, $aI\preceq\Sigma^{\prime}\preceq I.$ Using Lemma B.2, one can derive the following inequality:
@@
$(c)$ is obtained from the fact that $\tau\leq 2\sigma_{\min}(R\leq 2\|R+\gamma B^{\top}P_{K}B\|$, and $(d)$ follows from Lemma B.2. Meanwhile, observe from (50) that
@@
Finally, take $\zeta=\min\left\{\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S^{*}\|},\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\right\}$ and combine (54) and (55) to get
$C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)\leq-\zeta\left(C(K,\Sigma)-C(K^{*},\Sigma^{*})\right).$
C.3 Proof of Lemma 4.3
By (15) - (17),
@@
Appendix D Proofs in Section 5
D.1 Proof of Lemma 5.1
@@ Appendix D Proofs in Section 5
(of Lemma D.1). Let us first show the following equation holds for any $t=1,2,3,\cdots$ by induction:
@@
Lemma D.2 follows a direct calculation with triangle inequality.
Lemma D.3.
@@
(of Lemma D.3). To ease the exposition, let us denote $\mathcal{G}_{t}=\mathcal{G}^{{K_{1}}}_{t},\mathcal{G}^{\prime}_{t}=\mathcal{G}^{{K_{2}}}_{t},\mathcal{F}=\mathcal{F}_{K_{1}}$ and $\mathcal{F}^{\prime}=\mathcal{F}_{K_{2}}$. Then for any symmetric matrix $X\in\mathbb{R}^{n\times n}$ and $t=0,1,2,\cdots$,
@@
where $(a)$ follows from (62); $(b)$ follows from (59); and Lemma D.3. The constants $\xi_{\gamma,\rho},\zeta_{\gamma,\rho},\omega_{\gamma,\rho}$ are defined in (26). Finally, applying Lemma D.2 finishes the proof. ∎
D.4 Proof of Lemma 5.4
@@
Note that Lemma 5.3 in Appendix D.3 can be employed to derive a bound on $\|S^{(t+1)}-S^{*}\|$ in relation to $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|.$ In this section, we further establish this bound by deriving the bounds for $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|$ in terms of $\|K^{(t)}-K^{*}\|$ and $\|P_{K^{(t)}}-P_{K^{*}}\|$ (cf. Lemma D.4 and Lemma D.5). Additionally, the perturbation analysis for $P_{K}$ in Lemma D.7 demonstrates $\|P_{K^{(t)}}-P_{K^{*}}\|$ can be bounded by $\|{K^{(t)}}-{K^{*}}\|$, which completes the proof for Lemma 5.4.
We first establish the bound for the one-step update of $K$ and $\Sigma$ in Lemma D.4 and Lemma D.5 respectively.
Lemma D.4 (Bound of one-step update of $K$).
@@
$+\gamma\sigma_{\min}(R)\left(\|B\|\|A\|+\|B\|^{2}\|K^{(t)}\|\right)\cdot\|P_{K^{(t)}}-P_{K^{*}}\|.$
@@
(of Lemma D.4). Let $K^{\prime}$ denote $K^{(t+1)}$ and $K$ denote $K^{(t)}$ to ease the notation. Theorem 2.1 shows that for an optimal $K^{*}$,
@@
Combining (64) and (65) completes the proof. ∎
Lemma D.5 (Bound of one-step update of $\Sigma$).
@@
(of Lemma D.5). Observe from (IPO) and (6) that
@@
Lemma D.6 and Lemma D.7 perform perturbation analysis on $P_{K}$ and establish bounds for the difference in $P_{K}$ with respect to the perturbation in $K$. Consequently, both $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|$ (in Lemma D.4 and Lemma D.5) can be bounded in terms of $\|K^{(t)}-K^{*}\|$.
Lemma D.6 ($\mathcal{T}_{K}$ norm bound).
$\|\mathcal{T}_{K}\|\leq\frac{1}{\mu}\|S_{K,\Sigma}\|.$
@@
(of Lemma D.6). By Lemma 17 in [12], $\|\mathcal{T}_{K}\|\leq\frac{1}{\mu}\|\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])\|.$ Since $S_{K,\Sigma}\succeq\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])$, thus $\|\mathcal{T}_{K}\|\leq\frac{1}{\mu}\sigma_{max}\left(\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])\right)\leq\frac{1}{\mu}\|S_{K,\Sigma}\|.$
Lemma D.7 ($P_{K}$ perturbation).
Then for any $K_{1}$ and $K_{2}$ such that $\|A-BK_{1}\|,\|A-BK_{2}\|<\frac{1}{\sqrt{\gamma}}$,
$\|P_{K_{2}}-P_{K_{1}}\|$
$\leq\left(2\rho\xi_{\gamma\rho}\|B\|\left(\|Q\|+\|R\|\|K_{2}\|^{2}\right)+\frac{1}{\mu}\|S_{K_{1},\Sigma_{1}}\|\|R\|\cdot\left(\|K_{1}\|+\|K_{2}\|\right)\right)\left\|K_{1}-K_{2}\right\|.$
Proof.
(of Lemma D.7). Fix $K_{1}$ and $K_{2}$ such that $\|A-BK_{1}\|<\frac{1}{\sqrt{\gamma}}$ and $\|A-BK_{2}\|<\frac{1}{\sqrt{\gamma}}$. By (57) and (17), we have for any $K$,
@@
which immediately implies $P_{K}=\mathcal{T}_{K}(Q+K^{\top}RK).$ To bound the difference between $P_{K_{1}}$ and $P_{K_{2}}$, observe:
$\|P_{K_{2}}-P_{K_{1}}\|=\left\|\mathcal{T}_{{K_{2}}}\left(Q+K_{2}^{\top}R{K_{2}}\right)-\mathcal{T}_{K_{1}}\left(Q+{K_{1}}^{\top}R{K_{1}}\right)\right\|$
$\leq$ $\Big{\|}\mathcal{T}_{K_{2}}\left(Q+K_{2}^{\top}RK_{2}\right)-\mathcal{T}_{K_{1}}\left(Q+\left(K_{2}\right)^{\top}RK_{2}\right)$
$-\left(\mathcal{T}_{K_{1}}\left(Q+{K_{1}}^{\top}R{K_{1}}\right)-\mathcal{T}_{K_{1}}\left(Q+K_{2}^{\top}RK_{2}\right)\right)\Big{\|}$
$=$ $\left\|\mathcal{T}_{K_{2}}\left(Q+K_{2}^{\top}RK_{2}\right)-\mathcal{T}_{K_{1}}\left(Q+K_{2}^{\top}RK_{2}\right)-\mathcal{T}_{K_{1}}\circ\left({K_{1}}^{\top}R{K_{1}}-K_{2}^{\top}RK_{2}\right)\right\|$
$\leq$ $\left\|\mathcal{T}_{K_{2}}\left(Q+K_{2}^{\top}RK_{2}\right)-\mathcal{T}_{K_{1}}\left(Q+K_{2}^{\top}RK_{2}\right)\right\|+\|\mathcal{T}_{K_{1}}\|\|{K_{1}}^{\top}R{K_{1}}-K_{2}^{\top}RK_{2}\|.$ (66)
For the first term in (66), we can apply Lemma D.3 and Lemma D.2 to get
$\left\|\mathcal{T}_{K_{2}}\left(Q+K_{2}^{\top}RK_{2}\right)-\mathcal{T}_{K_{1}}\left(Q+K_{2}^{\top}RK_{2}\right)\right\|$ (67)
$\leq$ $\xi_{\gamma\rho}\|\mathcal{F}_{K_{1}}-\mathcal{F}_{K_{2}}\|\|Q+K_{2}^{\top}RK_{2}\|\leq 2\rho\xi_{\gamma\rho}\|B\|\left\|K_{1}-K_{2}\right\|\cdot\left(\|Q\|+\|R\|\|K_{2}\|^{2}\right).$
For the second term in (66), by Lemma D.6, we have
$\|\mathcal{T}_{K_{1}}\|\|{K_{1}}^{\top}R{K_{1}}-K_{2}^{\top}RK_{2}\|$ (68)
$=$ $\|\mathcal{T}_{K_{1}}\|\|{K_{1}}^{\top}R{K_{1}}-{K_{1}}^{\top}RK_{2}+{K_{1}}^{\top}RK_{2}-K_{2}^{\top}RK_{2}\|$
$\leq$ $\|\mathcal{T}_{K_{1}}\|\|R\|\|K_{2}-{K_{1}}\|\left(\|{K_{1}}\|+\|K_{2}\|\right)$
$\leq$ $\frac{\|S_{{K_{1}},\Sigma_{1}}\|}{\mu}\|R\|\|K_{2}-{K_{1}}\|\left(\|{K_{1}}\|+\|K_{2}\|\right).$
Plugging (67) and (68) in (66) completes the proof. ∎
@@
(of Lemma 5.4). Let $K^{\prime}$ denote $K^{(t+1)}$ and $K$ denote $K^{(t)}$ to ease the notation. Apply Lemma D.7 with $K$ and $K^{*}$ to get
@@
Apply Lemma D.4 with $\|K\|\leq\kappa$ and (69) to get
@@
Apply Lemma D.5 with (69) to get

### v3

@@ 8.2 Proofs in Section 3
$\overset{(b)}{\leq}(1-\gamma)\cdot\|(R+\gamma B^{\top}P_{K}B)^{-1}\|\cdot\|\nabla_{\Sigma}C(K,\Sigma)\|_{F}^{2}\leq\frac{(1-\gamma)\|\nabla_{\Sigma}C(K,\Sigma)\|_{F}^{2}}{\sigma_{\min}(R)},$
where $(a)$ follows from the first order concavity condition for $f_{K}$ and $(b)$ is from $0\prec\Sigma\preceq I.$
@@
8.2.3 Proof of Lemma 3.3
Lemma 8.2 shows that the cost objective is smooth in $\Sigma$ when utilizing entropy regularization, given that $\Sigma$ is bounded.
Lemma 8.2 (Smoothness of $f_{K}$ (16)).
@@
(of Lemma 8.2). Fix symmetric positive definite matrices $X$ and $Y$ satisfying $aI\preceq X\preceq I$ and $aI\preceq Y\preceq I$. Then $f_{K}$ being concave implies $f_{K}(X)-f_{K}(Y)+\operatorname{Tr}\left(\nabla f_{K}(X)^{\top}(Y-X)\right)\geq 0.$ To find an upper bound, observe that
@@
(of Lemma 3.3). The first equality immediately results from Lemma 8.1, by taking an expectation with respect to $x$ in (46). The last inequality follows directly from Lemma 8.2. ∎
8.3 Proofs in Section 4
8.3.1 Proofs of Lemma 4.3
@@ 8.3 Proofs in Section 4
8.3.2 Proof of Lemma 4.4
For ease of exposition, write $S=S_{K,\Sigma},$ $S^{\prime}=S_{K^{\prime},\Sigma^{\prime}}$, and $S^{*}=S_{K^{*},\Sigma^{*}}$. Let $M_{a}$ be defined in the same way as in Lemma 3.3. Let $f_{K}$ be defined as (16). Then Lemma 3.3 implies
$C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)$ $=\operatorname{Tr}\left(S^{\prime}(K^{\prime}-K)^{\top}(R+\gamma B^{\top}P_{K}B)(K^{\prime}-K)\right)$ (54)
@@
$\leq$ $4\eta_{1}^{2}\|R+\gamma B^{\top}P_{K}B\|\operatorname{Tr}\left(S^{\prime}E_{K}^{\top}E_{K}\right)-4\eta_{1}\operatorname{Tr}\left(S^{\prime}E_{K}^{\top}E_{K}\right)\overset{(a)}{\leq}-2\eta_{1}\operatorname{Tr}\left(S^{\prime}E_{K}^{\top}E_{K}\right)$
$\leq$ $-2\eta_{1}\mu\operatorname{Tr}\left(E_{K}^{\top}E_{K}\right)\overset{(b)}{\leq}-2\eta_{1}\mu\frac{\sigma_{\min}(R)}{\|S^{*}\|}\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right),$
where (a) follows from $\eta_{1}\leq\frac{1}{2\|R+\gamma B^{\top}P_{K}B\|}$ and (b) follows from (49).
By Lemma 4.3, $aI\preceq\Sigma^{\prime}\preceq I.$ Using Lemma 8.2, one can derive the following inequality:
@@
$(c)$ is obtained from the fact that $\tau\leq 2\sigma_{\min}(R)\leq 2\|R+\gamma B^{\top}P_{K}B\|$, and $(d)$ follows from Lemma 8.2. Meanwhile, observe from (50) that
@@
Finally, with $\zeta=\min\left\{\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S^{*}\|},\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\right\}$, plug (55) and (56) in (54) to get
$C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)$
$\leq-2\eta_{1}\mu\frac{\sigma_{\min}(R)}{\|S^{*}\|}\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)-\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\left(f_{K}(\Sigma)-f_{K}(\Sigma^{*})\right)$
$\leq-\zeta\left(\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)+f_{K}(\Sigma)-f_{K}(\Sigma^{*})\right)$
$\leq-\zeta(C(K,\Sigma)-C(K^{*},\Sigma^{*}))$
where the last inequality follows from (48). Adding $C(K,\Sigma)-C(K^{*},\Sigma^{*})$ to both sides of the above inequality finishes the proof.
8.3.3 Proof of Lemma 4.5
By (15) and (17),
@@
8.4 Proofs in Section 5
8.4.1 Proof of Lemma 5.2
@@ 8.4 Proofs in Section 5
(of Lemma 8.3). Let us first show the following equation holds for any $t=1,2,3,\cdots$ by induction:
@@
Lemma 8.4 follows a direct calculation with triangle inequality.
Lemma 8.5.
@@
(of Lemma 8.5). To ease the exposition, let us denote $\mathcal{G}_{t}=\mathcal{G}^{{K_{1}}}_{t},\mathcal{G}^{\prime}_{t}=\mathcal{G}^{{K_{2}}}_{t},\mathcal{F}=\mathcal{F}_{K_{1}}$ and $\mathcal{F}^{\prime}=\mathcal{F}_{K_{2}}$. Then for any symmetric matrix $X\in\mathbb{R}^{n\times n}$ and $t=0,1,2,\cdots$,
@@
where $(a)$ follows from (63); $(b)$ follows from (60); and Lemma 8.5. The constants $\xi_{\gamma,\rho},\zeta_{\gamma,\rho},\omega_{\gamma,\rho}$ are defined in (26). Finally, applying Lemma 8.4 finishes the proof. ∎
8.4.4 Proof of Lemma 5.6
@@
Note that Lemma 5.5 in Appendix 8.4.3 can be employed to derive a bound on $\|S^{(t+1)}-S^{*}\|$ in relation to $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|.$ In this section, we further establish this bound by deriving the bounds for $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|$ in terms of $\|K^{(t)}-K^{*}\|$ and $\|P_{K^{(t)}}-P_{K^{*}}\|$ (cf. Lemma 8.6 and Lemma 8.7). Additionally, the perturbation analysis for $P_{K}$ in Lemma 8.8 demonstrates $\|P_{K^{(t)}}-P_{K^{*}}\|$ can be bounded by $\|{K^{(t)}}-{K^{*}}\|$, which completes the proof for Lemma 5.6.
We first establish the bound for the one-step update of $K$ and $\Sigma$ in Lemma 8.6 and Lemma 8.7 respectively.
Lemma 8.6 (Bound of one-step update of $K$).
@@
$+\gamma\sigma_{\min}(R)\left(\|B\|\|A\|+\|B\|^{2}\kappa\right)\cdot\|P_{K^{(t)}}-P_{K^{*}}\|.$
@@
(of Lemma 8.6). Let $K^{\prime}$ denote $K^{(t+1)}$ and $K$ denote $K^{(t)}$ to ease the notation. Theorem 2.1 shows that for an optimal $K^{*}$,
@@
Combining (65) and (66), then using $\kappa\geq\|K\|$ for any $K\in\Omega$ completes the proof. ∎
Lemma 8.7 (Bound of one-step update of $\Sigma$).
@@
(of Lemma 8.7). Observe from (IPO) and (6) that
@@
Lemma 8.8 perform perturbation analysis on $P_{K}$ and establish bounds for the difference in $P_{K}$ with respect to the perturbation in $K$. Consequently, both $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|$ (in Lemma 8.6 and Lemma 8.7) can be bounded in terms of $\|K^{(t)}-K^{*}\|$.
Lemma 8.8 ($P_{K}$ perturbation).
For any $K\in\Omega,$ with $c$ defined in (27), $\|P_{K}-P_{K^{*}}\|\leq c\|K-K^{*}\|.$
@@
(of Lemma 8.8). Fix $K\in\Omega.$ By (58) and (17),
@@
which immediately implies $P_{K}=\mathcal{T}_{K}(Q+K^{\top}RK).$ Similarly $P_{K^{*}}=\mathcal{T}_{K^{*}}(Q+{K^{*}}^{\top}R{K^{*}}).$ To bound the difference between $P_{K}$ and $P_{K^{*}}$, observe that,
$\|P_{K}-P_{K^{*}}\|\leq\left\|\mathcal{T}_{K}\left(Q+K^{\top}RK\right)-\mathcal{T}_{K^{*}}\left(Q+K^{\top}RK\right)\right\|+\|\mathcal{T}_{K^{*}}\|\|{K^{*}}^{\top}R{K^{*}}-K^{\top}RK\|.$ (67)
For the first term in (67), we can apply Lemma 8.5 and Lemma 8.4 to get
$\left\|\mathcal{T}_{K}\left(Q+K^{\top}RK\right)-\mathcal{T}_{K^{*}}\left(Q+K^{\top}RK\right)\right\|$ (68)
$\leq$ $\xi_{\gamma\rho}\|\mathcal{F}_{K^{*}}-\mathcal{F}_{K}\|\|Q+K^{\top}RK\|\leq 2\rho\xi_{\gamma\rho}\|B\|\left\|K^{*}-K\right\|\cdot\left(\|Q\|+\|R\|\|K\|^{2}\right).$
For the second term in (67), note that by Lemma 17 in [12], $\|\mathcal{T}_{K}\|\leq\frac{1}{\mu}\|\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])\|.$ Since $S_{K,\Sigma}\succeq\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])$, thus $\|\mathcal{T}_{K}\|\leq\frac{1}{\mu}\sigma_{max}\left(\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])\right)\leq\frac{1}{\mu}\|S_{K,\Sigma}\|.$ Then we have
$\|\mathcal{T}_{K^{*}}\|\|{K^{*}}^{\top}R{K^{*}}-K^{\top}RK\|$ (69)
$=$ $\|\mathcal{T}_{K^{*}}\|\|{K^{*}}^{\top}R{K^{*}}-{K^{*}}^{\top}RK+{K^{*}}^{\top}RK-K^{\top}RK\|$
$\leq$ $\|\mathcal{T}_{K^{*}}\|\|R\|\|K-{K^{*}}\|\left(\|{K^{*}}\|+\|K\|\right)$
$\leq$ $\frac{\|S_{{K^{*}},\Sigma^{*}}\|}{\mu}\|R\|\|K-{K^{*}}\|\left(\|{K^{*}}\|+\|K\|\right).$
Plugging (68), (69), and (27) in (67) completes the proof. ∎
@@
(of Lemma 5.6). Let $K^{\prime}$ denote $K^{(t+1)}$ and $K$ denote $K^{(t)}$ to ease the notation. Apply Lemma 8.8 with $K$ and $K^{*}$ to get
@@
Apply Lemma 8.6 with $\|K\|\leq\kappa$ and (70) to get
@@
Apply Lemma 8.7 with (70) to get